### v1

# Sharp large deviation estimates for heavy-tailed extrema

Abstract.
We establish sharp large deviation asymptotics for the maximum order statistic of independent and identically distributed heavy-tailed random variables, valid for all Borel subsets of the right tail. This result yields exact decay rates for exceedance probabilities at thresholds that grow faster than the natural extreme-value scaling. As an application, we derive the polynomial rate of decay of ruin probabilities in insurance portfolios where insolvency is driven by a single extreme claim.
Key words: Heavy-tailed distributions; extreme value theory; large deviations; ruin probabilities; solvency risk.
AMS 2020 Subject Classification: 60G70, 60F10, 62G32, 62P05
The author acknowledges the partial support of the Ministerio de Ciencia e Innovación of Spain in the project PID2022-137396NB-I00, funded by MICIU/AEI/10.13039/501100011033 and by ’ERDF A way of making Europe’.
1. Main result
The present work establishes a new large deviation principle in the context of extreme value theory and analyzes its implications for the asymptotic behavior of extreme insurance claims under heavy-tailed loss distributions. Let $(X_{n})_{n\in\mathbb{N}}$ be independent and identically distributed random variables. We assume that the distribution has a heavy right tail, in the sense that its survival function $\bar{F}$ satisfies
$\bar{F}(x):=\mathbb{P}(X_{1}>x)=x^{-\alpha}L(x),\qquad x>x_{0},$
for some $x_{0}>0$, where $\alpha>0$ and $L$ is a slowly varying function, i.e.,
$\lim_{x\to\infty}\frac{L(tx)}{L(x)}=1,\qquad\text{for all }t>0.$
This assumption covers a wide range of claim size distributions commonly used in actuarial science and finance, including Pareto, Student-$t$, Burr, stable distributions with index $\alpha<2$, and Loggamma distributions; see Embrechts et al. [4].
Let $X_{(n)}:=\max_{1\leq i\leq n}X_{i}$ denote the maximum order statistic, which naturally arises in actuarial applications such as the modeling of the largest individual claim, extreme portfolio losses, or catastrophic risk over a fixed observation period.
A classical result in extreme value theory asserts that there exists a sequence $(a_{n})_{n\in\mathbb{N}}$ of positive normalizing constants such that
$\frac{X_{(n)}}{a_{n}}\;\xrightarrow{d}\;\Phi_{\alpha},$
where $\Phi_{\alpha}(x)=\exp(-x^{-\alpha})$, $x>0$, is the Fréchet distribution with shape parameter $\alpha>0$. A standard choice for the normalizing sequence $a_{n}$ is given by
$a_{n}:=F^{\leftarrow}\!\left(1-\frac{1}{n}\right),\qquad n\in\mathbb{N},$ (1.1)
where $F^{\leftarrow}(u):=\inf\{x\in\mathbb{R}:F(x)\geq u\}$ denotes the left-inverse of the distribution function $F$ of $X_{1}$. This result traces back to the pioneering work of Fisher and Tippett [5] and Gnedenko [6]; see also [4, 8] for comprehensive treatments of heavy-tailed distributions and regular variation.
While this result characterizes the typical scale of the maximum order statistic through weak convergence, it does not yield information on the rate of decay of tail probabilities of the form
$\mathbb{P}(X_{(n)}>c_{n}a_{n}),$
when the threshold sequence $(c_{n})_{n\in\mathbb{N}}$ increases with $n$. Such probabilities correspond to extreme scenarios in which the maximum order statistic exceeds its natural extreme-value scale by a growing factor. We provide further illustration of this topic and its relevance for capital assessment in insurance in Section 2.
To address this issue, we adopt a large deviation approach (see [3]) tailored to the maximum order statistic. Specifically, we investigate the large deviations of the rescaled sequence
$Z_{n}=\left(\frac{X_{(n)}}{a_{n}}\right)^{\frac{\alpha}{\log n}},$ (1.2)
where $a_{n}$ is defined by (1.1). We further assume that $X_{1}$ admits a density $f$ which is non-increasing for sufficiently large values of $x$ (this assumption can actually be relaxed; see Remark 3.5 below).
Our main result establishes a sharp large deviation principle for the right tail of the sequence $(Z_{n})_{n\in\mathbb{N}}$.
Theorem 1.1.
Let $(Z_{n})_{n\in\mathbb{N}}$ be defined by (1.2) under the assumptions above. Then, for every Borel set $A\subset[1,\infty)$,
$\lim_{n\to\infty}\frac{1}{\log n}\,\log\mathbb{P}(Z_{n}\in A)=-{\rm ess.inf}_{x\in A}\log x.$ (1.3)
Theorem 1.1 shows that the sequence $(Z_{n})_{n\in\mathbb{N}}$ satisfies a large deviation principle with speed $\log n$ and rate function $I(x)=\log x$ on $1,\infty)$. The large deviation principle is sharp in the sense that it yields an exact limit for the logarithmic asymptotics of tail probabilities, rather than upper and lower bounds as in the classical formulation of large deviation theory; see, for example, [3]. Moreover, the explicit form of the rate function implies a universal decay behavior that depends only on the tail index $\alpha$ and is insensitive to the slowly varying component of the underlying distribution.
The remainder of the paper is organized as follows. In Section 2, we study the implications of this result for estimating the decay of ruin probabilities for maximum claim portfolios. Section 3 is devoted to the proof of Theorem 1.1. The paper ends with a conclusion section.
2. Decay of ruin probabilities for maximum claim portfolios
In solvency regulation and risk management, a key objective is to understand how ruin probabilities behave when capital requirements are increased to protect against extreme but plausible losses. This is particularly relevant for portfolios exposed to heavy-tailed risks, where total losses may be dominated by a single exceptionally large claim, as is commonly observed in catastrophe, liability, and operational risk insurance; see, for example, [4, 7].
Consider an insurance portfolio consisting of $n$ independent policies, and let $X_{1},\ldots,X_{n}$ denote the corresponding claim sizes. The insurer charges a premium $\pi_{n}$ per policy, so that the total available capital equals $n\pi_{n}$. We focus on a ruin mechanism driven by a single catastrophic claim and define ruin to occur if the largest claim exceeds the total premium income. The corresponding ruin probability is
$\mathrm{RP}_{n}=\mathbb{P}(X_{(n)}>n\pi_{n}),\qquad X_{(n)}:=\max_{1\leq i\leq n}X_{i}.$
If the premium is chosen at the classical extreme value level $\pi_{n}=a_{n}n^{-1}$, see (1.1), then the Fréchet limit implies
$\mathrm{RP}_{n}\longrightarrow 1-\Phi_{\alpha}(1)>0,$
so that the probability of ruin does not vanish as the portfolio size grows. This regime is therefore insufficient for ensuring asymptotic solvency.
To study the effect of more conservative capital requirements, we consider a premium of the form
$\pi_{n}=a_{n}n^{\beta-1},\qquad\beta>0,$
which corresponds to a polynomial increase in capital relative to the standard extreme value scaling. In this case, the ruin probability satisfies
$\mathrm{RP}_{n}=\mathbb{P}(X_{(n)}>a_{n}n^{\beta})\longrightarrow 0,\qquad n\to\infty,$
ensuring asymptotic safety. However, classical extreme value theory does not quantify how fast this probability converges to zero. In contrast, using Theorem 1.1, we obtain a precise asymptotic rate. Indeed, noting that
$X_{(n)}>n\pi_{n}\quad\Longleftrightarrow\quad Z_{n}>e^{\alpha\beta},$
the large deviation principle (1.3) yields
$\lim_{n\to\infty}\frac{1}{\log n}\log\mathrm{RP}_{n}=-\alpha\beta.$ (2.1)
Equivalently,
$\mathrm{RP}_{n}=n^{-\alpha\beta+o(1)},\qquad n\to\infty.$ (2.2)
Expression (2.2) shows that ruin probabilities decay at a polynomial rate in the portfolio size. The tail index $\alpha$ determines the sensitivity of solvency to capital stress, while the parameter $\beta$ controls the degree of conservatism in premium loading.
From an actuarial perspective, this result offers a tool for capital assessment, since the large deviation approach yields the correct asymptotic order of ruin probabilities in extreme regimes.
3. Proof of the main result
In this section, we provide the proof of the main result Theorem 1.1. Let $(X_{n})_{n\geq 1}$ be i.i.d. random variables satisfying the assumptions in Section 1, and let $X_{(n)}=\max_{1\leq i\leq n}X_{i}$. Let $Z_{n}$ be defined by (1.2). Introduce the notation
$t_{n}(x)=a_{n}x^{\frac{\log n}{\alpha}},\qquad I(x)=\log x.$
Let $G_{n}$ and $g_{n}$ denote the distribution function and density of $Z_{n}$, respectively. A direct computation yields
$G_{n}(x)=F^{n}\bigl(t_{n}(x)\bigr),\qquad g_{n}(x)=n\,a_{n}\,\frac{\alpha}{\log n}\,x^{\frac{\log n}{\alpha}-1}\,F^{\,n-1}\!\bigl(t_{n}(x)\bigr)\,f\!\bigl(t_{n}(x)\bigr).$ (3.1)
The following lemma describes the asymptotic behaviour of the normalizing sequence $(a_{n})$.
Lemma 3.1.
It holds that
$\lim_{n\to\infty}\frac{\log a_{n}}{\log n}=\frac{1}{\alpha}.$ (3.2)
In particular,
$a_{n}=n^{\frac{1}{\alpha}+o(1)}\qquad\text{as }n\to\infty.$ (3.3)
Proof.
By [8, Proposition 0.8(ii)], the tail distribution satisfies
$\lim_{x\to\infty}\frac{\log\bar{F}(x)}{\log x}=-\alpha.$ (3.4)
We first note that $a_{n}\to\infty$. Indeed, if $(a_{n})$ were bounded, then $\bar{F}(x)=0$ for all sufficiently large $x$, implying that $X_{1}$ is almost surely bounded, which contradicts (3.4).
Fix $\varepsilon\in(0,1)$. By (3.4), there exists $x_{\varepsilon}>0$ such that for all $x\geq x_{\varepsilon}$,
$x^{-(\alpha+\varepsilon)}\;\leq\;\bar{F}(x)\;\leq\;x^{-(\alpha-\varepsilon)}.$ (3.5)
For $n$ large enough, $a_{n}-\varepsilon\geq x_{\varepsilon}$. By definition (1.1) of $a_{n}$,
$\bar{F}(a_{n})\leq\frac{1}{n}.$ (3.6)
Using the lower bound in (3.5) together with (3.6), we obtain
$a_{n}^{-(\alpha+\varepsilon)}\leq\frac{1}{n},$
which implies
$a_{n}\geq n^{\frac{1}{\alpha+\varepsilon}}.$ (3.7)
Moreover, by the definition of $a_{n}$ as an infimum, we have for every fixed $\varepsilon>0$,
$\bar{F}(a_{n}-\varepsilon)\geq\frac{1}{n},$ (3.8)
for all $n$ sufficiently large.
Applying the upper bound in (3.5) at $x=a_{n}-\varepsilon$, we obtain
$\frac{1}{n}\;\leq\;(a_{n}-\varepsilon)^{-(\alpha-\varepsilon)}.$
Hence,
$a_{n}\;\leq\;\varepsilon+n^{\frac{1}{\alpha-\varepsilon}}.$ (3.9)
Combining (3.7) and (3.9), we obtain
$\frac{1}{\alpha+\varepsilon}\;\leq\;\frac{\log a_{n}}{\log n}\;\leq\;\frac{1}{\alpha-\varepsilon}+\frac{\log\!\left(1+\varepsilon n^{-\frac{1}{\alpha-\varepsilon}}\right)}{\log n}.$
Letting $n\to\infty$ yields
$\frac{1}{\alpha+\varepsilon}\leq\liminf_{n\to\infty}\frac{\log a_{n}}{\log n}\leq\limsup_{n\to\infty}\frac{\log a_{n}}{\log n}\leq\frac{1}{\alpha-\varepsilon}.$
Since $\varepsilon>0$ is arbitrary, letting $\varepsilon\downarrow 0$ proves (3.2), and (3.3) follows immediately. ∎
The following result is classical in the theory of regularly varying functions and is commonly referred to as the Potter bounds; see, for instance, [8, Proposition 0.8(ii), p. 22] or [1, Theorem 1.5.6].
Lemma 3.2.
For every $\varepsilon>0$ there exists $t_{0}>0$ such that
$(1-\varepsilon)\,x^{-\alpha-\varepsilon}\;\leq\;\frac{\bar{F}(tx)}{\bar{F}(t)}\;\leq\;(1+\varepsilon)\,x^{-\alpha+\varepsilon},$
for all $t\geq t_{0}$ and all $x\geq x_{0}$.
We first establish an upper bound for the right tail probabilities, which corresponds to the large-deviation upper bound in Theorem 1.1.
Lemma 3.3.
For every $x\geq 1$,
$\limsup_{n\to\infty}\frac{1}{\log n}\log\mathbb{P}(Z_{n}>x)\leq-I(x).$
Recall that $I(x)=\log x$.
Proof.
Fix $x\geq 1$. Using the elementary identity
$1-u^{n}=(1+u+\cdots+u^{n-1})(1-u),\qquad u\in[0,1],$
we obtain the inequality
$1-u^{n}\leq n(1-u),\qquad u\in[0,1].$
Applying this with $u=F(t_{n}(x))$, we obtain
$\mathbb{P}(Z_{n}>x)=1-F(t_{n}(x))^{n}\leq n\,\bar{F}(t_{n}(x)).$
Since $x\geq 1$, we have $t_{n}(x)\geq a_{n}\to\infty$ as $n\to\infty$. Fix $\varepsilon>0$. By the Potter bounds (Lemma 3.2), for all $n$ sufficiently large,
$\bar{F}(t_{n}(x))\leq(1+\varepsilon)\,\bar{F}(a_{n})\,x^{\frac{\log n}{\alpha}(-\alpha+\varepsilon)}.$
Since $\bar{F}(a_{n})\leq 1/n$, we deduce
$\mathbb{P}(Z_{n}>x)\leq(1+\varepsilon)\,x^{\frac{\log n}{\alpha}(-\alpha+\varepsilon)}.$
Taking logarithms, dividing by $\log n$ , and taking the limit superior yield
$\limsup_{n\to\infty}\frac{1}{\log n}\log\mathbb{P}(Z_{n}>x)\leq\left(-1+\frac{\varepsilon}{\alpha}\right)\log x.$
Letting $\varepsilon\downarrow 0$ completes the proof. ∎
Lemma 3.4.
The density $g_{n}$ of $Z_{n}$ satisfies
$\lim_{n\to\infty}\frac{1}{\log n}\log g_{n}(x)=-I(x),$
with convergence uniform for $x\in[1,M]$, for any fixed $M>1$.
Proof.
Recall that the density $f$ of $X_{1}$ is assumed to be non-increasing for $x$ large enough. Under this assumption, the Von Mises condition holds:
$\lim_{x\to\infty}\frac{xf(x)}{\bar{F}(x)}=\alpha,$ (3.10)
see [8, Proposition 1.15(b)]. This relation will be used below.
From (3.1), we write
$\frac{1}{\log n}\log g_{n}(x)=T_{1}(n)+T_{2}(n,x)+T_{3}(n,x)+T_{4}(n,x),$
where
$T_{1}(n)=\frac{1}{\log n}\log\!\left(n\,a_{n}\,\frac{\alpha}{\log n}\right),\quad T_{2}(n,x)=\left(\frac{1}{\alpha}-\frac{1}{\log n}\right)\log x,$
$T_{3}(n,x)=\frac{n-1}{\log n}\log F\!\bigl(t_{n}(x)\bigr),\quad T_{4}(n,x)=\frac{1}{\log n}\log f\!\bigl(t_{n}(x)\bigr).$
By Lemma 3.1, $\log a_{n}=(1/\alpha+o(1))\log n$, which immediately yields
$\lim_{n\to\infty}T_{1}(n)=1+\frac{1}{\alpha}.$
In addition,
$\lim_{n\to\infty}T_{2}(n,x)=\frac{1}{\alpha}\log x,$
uniformly for $x\in[1,M]$.
Since $x\geq 1$, we have $t_{n}(x)\geq a_{n}\to\infty$. Moreover,
$0\leq-\log F(t_{n}(x))\leq-\log F(a_{n}).$
For $u$ sufficiently close to $1$, the inequality $0\leq-\log u\leq 2(1-u)$ holds, and hence
$\log F(a_{n})\leq 2\bar{F}(a_{n})\leq\frac{2}{n},$
where we have used that $\bar{F}(a_{n})\leq 1/n$. It follows that
$|T_{3}(n,x)|\leq\frac{n-1}{\log n}\frac{2}{n},$
which converges to 0 as $n\to\infty$, uniformly for $x\in[1,M]$.
We now study the fourth term
$T_{4}(n,x)=\frac{1}{\log n}\log f\bigl(t_{n}(x)\bigr).$
Since the density $f$ is eventually non-increasing, the Von Mises condition (3.10) holds. Consequently, for $n$ large enough and all $x\in[1,M]$, we have
$(\alpha-1)\frac{\bar{F}(t_{n}(x))}{t_{n}(x)}\leq f(t_{n}(x))\leq(\alpha+1)\frac{\bar{F}(t_{n}(x))}{t_{n}(x)}.$ (3.11)
Using Potter bounds iteratively, we may construct a sequence $\varepsilon_{n}\downarrow 0$ such that, for all $n$ sufficiently large and all $x\in[1,M]$,
$\frac{1-\varepsilon_{n}}{n}\,x^{-\frac{\log n}{\alpha}(\alpha+\varepsilon_{n})}\leq\bar{F}(t_{n}(x))\leq\frac{1+\varepsilon_{n}}{n}\,x^{-\frac{\log n}{\alpha}(\alpha-\varepsilon_{n})}.$ (3.12)
Combining the upper bounds in (3.11) and (3.12), we obtain
$\frac{1}{\log n}\log f\bigl(t_{n}(x)\bigr)\leq\frac{1}{\log n}\log\!\left(\frac{\alpha+1}{t_{n}(x)}\right)+\frac{1}{\log n}\log\!\left(\frac{1+\varepsilon_{n}}{n}\right)-\left(1-\frac{\varepsilon_{n}}{\alpha}\right)\log x.$
Using Lemma 3.1, we deduce that, uniformly for $x\in[1,M]$,
$\frac{1}{\log n}\log f\bigl(t_{n}(x)\bigr)\leq-\left(1+\frac{1}{\alpha}\right)-\left(1+\frac{1}{\alpha}\right)\log x+\frac{\varepsilon_{n}}{\alpha}\log M+o(1).$
Similarly, combining the lower bounds in (3.11) and (3.12), we obtain
$\frac{1}{\log n}\log f\bigl(t_{n}(x)\bigr)\geq-\left(1+\frac{1}{\alpha}\right)-\left(1+\frac{1}{\alpha}\right)\log x-\frac{\varepsilon_{n}}{\alpha}\log M+o(1).$
Together, these inequalities imply that
$\left|\frac{1}{\log n}\log f\bigl(t_{n}(x)\bigr)+\left(1+\frac{1}{\alpha}\right)+\left(1+\frac{1}{\alpha}\right)\log x\right|\leq\frac{\varepsilon_{n}}{\alpha}\log M+o(1),$
uniformly for $x\in[1,M]$. Since $\varepsilon_{n}\to 0$, we conclude that
$\lim_{n\to\infty}T_{4}(n,x)=-\left(1+\frac{1}{\alpha}\right)-\left(1+\frac{1}{\alpha}\right)\log x,$
uniformly for $x\in[1,M]$.
Collecting the limits of $T_{1}$, $T_{2}$, $T_{3}$, and $T_{4}$, we conclude that
$\lim_{n\to\infty}\frac{1}{\log n}\log g_{n}(x)=-\log x=-I(x),$
uniformly for $x\in[1,M]$. This completes the proof. ∎
We finally prove Theorem 1.1.
Proof.
Fix a Borel set $A\subset[1,\infty)$ with positive Lebesgue measure; otherwise the statement is trivial. Set
$I_{A}:={\rm ess.inf}_{x\in A}I(x).$
Fix $\varepsilon>0$ and define
$A_{\varepsilon}:=\{x\in A:\ I(x)<I_{A}+\varepsilon\}.$
By definition of the essential infimum, $A_{\varepsilon}$ has positive Lebesgue measure. Choose $M>0$ large enough so that $A_{\varepsilon}\cap[0,M]$ also has positive Lebesgue measure.
By the uniform convergence established in Lemma 3.4, there exists $n_{0}\in\mathbb{N}$ such that
$n^{-I(x)-\varepsilon}\leq g_{n}(x)\leq n^{-I(x)+\varepsilon},\qquad\text{for all }x\in[0,M],\ n\geq n_{0}.$ (3.13)
We now derive matching lower and upper bounds for the logarithmic asymptotics.
First, for the lower bound, we obtain
$\liminf_{n\to\infty}\frac{1}{\log n}\log\mathbb{P}(Z_{n}\in A)$ $\geq\liminf_{n\to\infty}\frac{1}{\log n}\log\mathbb{P}(Z_{n}\in A_{\varepsilon}\cap[0,M])$
$=\liminf_{n\to\infty}\frac{1}{\log n}\log\int_{A_{\varepsilon}\cap[0,M]}g_{n}(x)\,dx$
$\geq\liminf_{n\to\infty}\frac{1}{\log n}\log\int_{A_{\varepsilon}\cap[0,M]}n^{-I(x)-\varepsilon}\,dx$
$\geq\liminf_{n\to\infty}\frac{1}{\log n}\log\int_{A_{\varepsilon}\cap[0,M]}n^{-I_{A}-2\varepsilon}\,dx$
$=-I_{A}-2\varepsilon,$
where we used that $I(x)\leq I_{A}+\varepsilon$ on $A_{\varepsilon}$ and that the Lebesgue measure of $A_{\varepsilon}\cap[0,M]$ is positive.
Next, for the upper bound, we decompose
$\mathbb{P}(Z_{n}\in A)\leq\mathbb{P}(Z_{n}\in A\cap[0,M])+\mathbb{P}(Z_{n}>M)\leq 2\max\{\mathbb{P}(Z_{n}\in A\cap[0,M]),\mathbb{P}(Z_{n}>M)\}.$
Taking logarithms and dividing by $\log n$, we obtain
$\limsup_{n\to\infty}\frac{1}{\log n}\log\mathbb{P}(Z_{n}\in A)\leq\max\Bigg\{$ $\limsup_{n\to\infty}\frac{1}{\log n}\log\mathbb{P}(Z_{n}\in A\cap[0,M]),$
$\limsup_{n\to\infty}\frac{1}{\log n}\log\mathbb{P}(Z_{n}>M)\Bigg\}.$
By Lemma 3.3, we have
$\limsup_{n\to\infty}\frac{1}{\log n}\log\mathbb{P}(Z_{n}>M)\leq-I(M).$
On the other hand, using the upper bound in (3.13),
$\limsup_{n\to\infty}\frac{1}{\log n}\log\mathbb{P}(Z_{n}\in A\cap[0,M])$ $\leq\limsup_{n\to\infty}\frac{1}{\log n}\log\int_{A\cap[0,M]}n^{-I(x)+\varepsilon}\,dx$
$\leq-I_{A}+\varepsilon.$
Combining the lower and upper bounds yields
$I_{A}-2\varepsilon\leq\liminf_{n\to\infty}\frac{1}{\log n}\log\mathbb{P}(Z_{n}\in A)\leq\limsup_{n\to\infty}\frac{1}{\log n}\log\mathbb{P}(Z_{n}\in A)\leq\max\{-I_{A}+\varepsilon,\,-I(M)\}.$
Finally, letting $M\uparrow\infty$ and then $\varepsilon\downarrow 0$, we obtain
$\lim_{n\to\infty}\frac{1}{\log n}\log\mathbb{P}(Z_{n}\in A)=-I_{A},$
which completes the proof. ∎
We discuss the assumptions of Theorem 1.1.
Remark 3.5.
While the von Mises condition (3.10) holds whenever $f$ is eventually monotone, the proof of Theorem 1.1 only requires the weaker assumption that the ratio
$\left|\frac{xf(x)}{\bar{F}(x)}\right|$ (3.14)
is bounded for sufficiently large $x$.
By [8, Corollary 1.12], $\bar{F}$ has the representation
$\bar{F}(x)=c(x)\exp\left(-\int_{1}^{x}\frac{\alpha(t)}{t}dt\right).$
for $x\geq 1$, where $\lim_{x\to\infty}c(x)=c>0$ and $\lim_{x\to\infty}\alpha(x)=\alpha>0$.
By differentiating the above representation, we get
$\frac{xf(x)}{\bar{F}(x)}=\alpha(x)-\frac{xc^{\prime}(x)}{c(x)}.$
Since $c(x)$, $\alpha(x)$ converge to positive constants, they are bounded for large $x$. Consequently, to ensure the boundedness of (3.14), it suffices to assume that $|xc^{\prime}(x)|$ is bounded for $x$ large enough, or equivalently, $c^{\prime}(x)=O(1/x)$. So the monotonicity of $f$ (or the Von Mises condition) can be replaced by this condition in the proof of Theorem 1.1 to hold. Note that in the common case where $c(x)$ is constant for large $x$ (so $c^{\prime}(x)=0$), the ratio simply converges to $\alpha$, recovering the classical von Mises condition (3.10).
4. Conclusion
In this paper, we established a novel sharp large deviation principle for the maximum of i.i.d. heavy-tailed random variables. By introducing a logarithmic rescaling of the classical extreme-value normalization, we derived tail probability estimates that are valid on all Borel subsets of the extreme right tail. The resulting large deviation principle has speed $\log n$ and an explicit rate function, leading to polynomial decay of extreme tail probabilities in the sample size $n$.
From an actuarial and risk-management perspective, the derived asymptotics yield explicit decay rates for ruin probabilities in models where insolvency is driven by a single catastrophic claim. The resulting polynomial rates depend only on the tail index $\alpha$ and are insensitive to the slowly varying component of the claim size distribution, which is often difficult to estimate reliably in practice.
The analysis highlights the limitations of classical extreme value approximations when capital levels increase above their natural scaling. While the Fréchet limit accurately describes typical extremes, it does not capture the speed at which ruin probabilities vanish under increasingly conservative capital requirements. The large deviation framework developed in this paper fills this gap by providing asymptotic estimates that remain informative in rare-event regimes relevant for stress testing and capital adequacy assessment.